# ASVspoof 5: Design, Collection and Validation of Resources for Spoofing, Deepfake, and Adversarial Attack Detection Using Crowdsourced Speech

### v3

@@ Abstract
ASVspoof 5 is the fifth edition in a series of challenges which promote the study of speech spoofing and deepfake attacks as well as the design of detection solutions. We introduce the ASVspoof 5 database which is generated in crowdsourced fashion from data collected in diverse acoustic conditions (cf. studio-quality data for earlier ASVspoof databases) and from $\sim$2,000 speakers (cf. $\sim$100 earlier). The database contains attacks generated with 32 different algorithms, also crowdsourced, and optimised to varying degrees using new surrogate detection models. Among them are attacks generated with a mix of legacy and contemporary text-to-speech synthesis and voice conversion models, in addition to adversarial attacks which are incorporated for the first time. ASVspoof 5 protocols comprise seven speaker-disjoint partitions. They include two distinct partitions for the training of different sets of attack models, two more for the development and evaluation of surrogate detection models, and then three additional partitions which comprise the ASVspoof 5 training, development and evaluation sets. An auxiliary set of data collected from an additional 30k speakers can also be used to train speaker encoders for the implementation of attack algorithms. Also described herein is an experimental validation of the new ASVspoof 5 database using a set of automatic speaker verification and spoof/deepfake baseline detectors. With the exception of protocols and tools for the generation of spoofed/deepfake speech, the resources described in this paper, already used by participants of the ASVspoof 5 challenge in 2024, are now all freely available to the community.
@@ 1 Introduction
The potential threats and risks related to the misuse of deepfakes—synthetic or manipulated media generated with the aid of deep learning—are well acknowledged both by various professional communities (e.g. security, forensics, biometrics) and government bodies, as well as by the general public. Having emerged as a novel type of information security threat which covers multiple domains—news, social media, and communication to name a few—it is more timely than ever to develop proactive defences against deepfakes. The focus of this article is specifically speech deepfakes. Speech, as the primary means of human communication, carries not only the linguistic message but also paralinguistic (beyond language) information. This include voice timbre in addition to other personally identifiable attributes, representations of which can be estimated from just a few seconds of speech and then used to infer the speaker identity using automatic speaker verification (ASV) technology. The voice timbre can then be transplanted into synthetic or manipulated speech to impersonate a specific, target individual.
@@ 2.2 ASVspoof 5 database partitions
Figure 2: An illustration of the MLS English source database partitioning scheme and its use in generation of the ASVspoof 5 database, including training, development and evaluation sets illustrated in gray boxes. Numbers in the second row indicate the number of speakers in each partition. Those in rectangular blocks denote the number of utterances. The green box denotes the 8 TTS/VC systems used for generation of the ASVspoof 5 training set. The blue box denotes 8 TTS/VC systems used for generation of the development set whereas the red box denotes the 16 TTS/VC systems and adversarial attacks used for generation of the evaluation set. Solid black and coloured arrows denote use of data for the training of TTS/VC attacks or surrogate models. Dashed arrows denote the adaptation towards target speaker and attack generation using input text or source speech utterance.
@@ 3.1 Attack types
TTS-based attacks are performed using a synthetic utterance which contains the voice of a target speaker learned from one or more adaptation utterances and the linguistic content of a provided text input. As illustrated in Figure 4, TTS systems follow one of two general approaches, namely zero-shot voice cloning or few-shot speaker adaptive TTS.121212 Some legacy approaches such as unit-selection TTS are neither zero-shot nor few-shot in strict terms. To help clarify the concept, we consider the operation of the unit-selection TTS, which is used in the ASVspoof 5 database, as a special form of few-shot speaker-adaptive TTS. These systems synthesize speech by selecting and concatenating small speech segments called units. The acoustic decoders (i.e., the selection algorithms) extract units from the adaptation utterance(s). The resulting sequence of units is then simply concatenated by the ‘vocoder’. Readers are encouraged to refer to TTS literature for an in-depth explanation of legacy approaches [37]. A text encoder is used to transform the input text into features which (explicitly or implicitly) encode word pronunciation information and other suprasegmental attributes (e.g., pitch and rhythm). The acoustic decoder then produces acoustic features (e.g., mel-spectrograms), which are further transformed by the vocoder into a speech waveform. [^13] The modules are not necessarily implemented in the form of DNNs. They can instead be a set of linguistic rules or decision trees. So-called end-to-end TTS systems in the form of a single DNN can be divided conceptually into distinct blocks, each of which corresponds to one of the modules illustrated in Figure 4. The two approaches illustrated in Figure 4 differ in how the target speaker voice is learned. In the case of zero-shot voice cloning (top-middle), a target speaker embedding is extracted from the adaptation utterances by a speaker encoder. The acoustic decoder, and optionally also the text encoder and vocoder, are then conditioned upon the target speaker embedding. Since they use target speaker embeddings, zero-shot voice cloning TTS systems are pre-trained using data collected from a large number of speakers, but are not fine-tuned at the level of the target speaker. This sets zero-shot voice cloning systems apart from few-shot speaker adaptive TTS systems (bottom-middle) which are fined-tuned to the voice of the target speaker using the set of adaptation utterances and their transcriptions.
@@
VC-based attacks are implemented using an input utterance containing the voice of a non-target speaker in place of input text, but again using one or more adaptation utterances which contain the voice of a target speaker. The voice in the input utterance is converted to that of the target speaker, whereas the linguistic content of the former is preserved. Approaches to VC can be categorized similarly to the TTS systems illustrated in Figure 4, though with input utterances in place of input text and with a waveform encoder in place of the text encoder. VC acoustic decoders typically use similar statistical models or deep neural networks (DNNs) as TTS, but with necessary modifications to accommodate input features of different dimension.
@@
A28 Zero-shot TTS TF encoder Normalizing flow H/ASP HiFi-GAN YourTTS, pretrained
A29 Zero-shot TTS GPT-2 - H/ASP HiFi-GAN XTTS [55], pretrained
@@ 3.2 Training set
A01: a neural TTS system based on Glow-TTS [43]. At the core is an acoustic decoder which generates mel-spectrograms via a normalizing-flow-based deep generative model named Glow [56]. The acoustic decoder is conditioned on latent linguistic features which are extracted from the input text by a text encoder using the Transformer encoder architecture [57]. The vocoder is a generative adversarial network (GAN)-based DNN named HiFi-GAN [58]. The speaker encoder is an ECAPA-TDNN [59]. Glow-TTS is open-sourced and is widely used within the TTS community.
@@
A04: a neural TTS system based on Grad-TTS [44], featuring a diffusion-based acoustic decoder. The acoustic decoder generates a mel-spectrogram via an iterative reverse diffusion process [62]. From latent linguistic features generated by the text encoder, the acoustic decoder with a U-Net architecture [63] is used to predict a ‘residual’ (or, more specifically, the gradient field which maximizes the probability of data to be generated). The residual is then summed with the latent features, the output of which is again fed to the acoustic decoder to produce another residual. The process is repeated ten times, and the output of the last iteration is used as a generated mel-spectrogram. The text encoder is a Transformer encoder with text and speaker embedding inputs. The vocoder is a HiFi-GAN. The implementation is based on the code released by the original paper, but speaker embeddings are extracted using an ECAPA-TDNN speaker encoder.
@@
A08: a variational inference with adversarial learning for end-to-end text-to-Speech (VITS) [46] system which uses a Transformer-based text encoder and a HiFi-GAN vocoder. The acoustic decoder is based on normalizing flow, similar to A01-A03. However, the text encoder, the acoustic decoder, and the vocoder are jointly optimised by maximizing a variational evidence lower bound [46]. Speaker embeddings are x-vectors [64]. VITS produces especially high quality synthetic speech and is readily accessible with the open-sourced toolkit ESPNet [65].
@@ 3.3 Development set
A11: a zero-shot TTS system based on Tacotron 2 [48]. The speaker encoder is a three-layer recurrent neural network (RNN) with long-short-term-memory (LSTM) units [71], trained for a speaker diarization task using generalized end-to-end (GE2E) loss [72]. The text encoder comprises convolution and recurrent layers. The acoustic decoder uses a location-sensitive attention block [73] and an autoregressive (AR) DNN [48] to generate a mel-spectrogram which is then converted into a speech waveform using the diffusion-based WaveGrad [74] vocoder. Tacotron 2 has been reported to produce speech of near-to-natural quality for single-speaker, neutral style TTS [48].
@@
A13: a zero-shot VC (ZSVC) system based on the StarGAN-ZSVC model [49]. The waveform encoder extracts a mel-spectrogram from the input utterance. The StarGAN [75] acoustic decoder then converts the mel-spectrogram so that it encodes the same linguistic content but the voice of the target speaker. The speaker embedding is extracted using a stack of recurrent layers using gated recurrent units (GRUs) [76]. The vocoder is WaveGlow [77], a normalizing-flow-based waveform generation model. StarGANs are known to perform well in voice conversion tasks  [78] and the StarGAN-ZSVC model is available as open-source [49].
A14: a zero-shot neural TTS system built upon YourTTS [50], which is itself based on VITS [46]. A14 is an variant of A08. Advances include the higher number of Transformer blocks used in the text encoder and use of a RawNet3 speaker encoder [79].
@@ 3.4 Evaluation set
A24: a zero-shot VC system. An RNN-CNN waveform encoder is used to extract a phonetic posteriorgram [97, 83], a latent representation of mostly speaker-independent phonetic information. The acoustic decoder, also a CNN, combines the posteriorgam with ECAPA-TDNN speaker embeddings to predict target-speaker-dependent latent features. A HiFi-GAN vocoder is then used to produce speech in the voice of the target speaker. All components are trained in an end-to-end manner. A24 is similar to A16 in that input speech is disentangled into speaker-dependent and speaker-independent features. A key difference to A16 is use of GAN-based, end-to-end training.
@@
A29: another pre-trained, zero-shot TTS system, named XTTS [55]. Designed for multi-lingual TTS and trained using a mix of four datasets containing speech in 16 languages, it uses the GPT-2 text encoder architecture [101], a HiFi-GAN vocoder, and an H/ASP speaker encoder (same as for A28). The XTTS stystem is also publicly available in the Coqui toolkit.
@@ 4.1 Encoding and compression
Condition C00 serves as a reference for which there is no encoding or compression. For conditions C00-C07, all data is sampled at a rate of 16 kHz. In contrast, conditions C08-C11 correspond to a narrow band setting under which all data is down-sampled to 8 kHz. Conditions C01-C11 all involve the application of lossy encoding/compression algorithms at the bitrates indicated in Table 3. C04 and C07 use the recent, deep-learning-based Encodec [102] scheme. For C07, mp3 encoding is applied prior to Encodec to simulate a channel with double compression. Finally, C11 is an experimental setting in which utterances are linearly convolved with the pre-computed, non-linear responses of a set of eight end-to-end calling pipelines, namely calls made from a device to a call center platform over a public switched telephone network (PSTN). Responses were estimated using the syncronized swept sine approach described in [103] using one of six different calling devices and one of four different audio injection methods. Specific configurations for C11 are detailed in A.
@@ 4.2 Shortcut artefacts
We created a post-processing pipeline to detect and reduce the impact of potential shortcut artefacts. It was applied to the ASVspoof 5 evaluation and development sets, but not to the training set. This choice encourages database users to explore techniques which are able to extract the most semantically relevant cues while also avoiding the pitfalls of shortcut learning. Our analysis revealed five potential shortcut artefacts: the peak waveform amplitude; the durations of the leading and trailing non-speech segments [13, 14]; the duration and energy of the whole utterance. The non-speech and speech segments of the utterance in the ASVspoof 5 evaluation set are annotated using the Whisper system [105]. Figure 5 shows the distribution of each shortcut artefact for bona fide utterances (black profiles) and spoofed/deepfake utterances (gray profiles) before post-processing. Distributions are shown in the first three rows for utterances generated using A17 (TTS), A21 (TTS), and A25 (VC) attacks. They show that, for A17 and A25 attacks, the peak amplitude is rescaled to 1.0, a distinct difference to the distribution for bona fide utterances. Whereas there is little difference between the distributions of leading non-speech segment durations for bona fide and spoofed/deepfake utterances for A25, they tend to be of shorter duration for A21. For A17, the duration of leading non-speech segments is uniformly distributed, even if the range is similar to that of bona fide utterances. Even so, the range in the duration of trailing non-speech segments for A17 differs to that for bona fide utterances. The duration of trailing non-speech segments for A21 is generally again shorter than those for bona fide utterances. The distribution of total duration for A25 overlaps with that of bona fide data. This is because A25, a VC system, does not alter the total duration of the input utterance. In the case of TTS-based attacks A17 and A21, the total duration is controlled by the acoustic model, hence the difference between distributions for bona fide and spoofed/deepfake utterances in these cases. There are also distinct differences in distributions of average energy for all three attacks.
@@
Figure 5: Distributions of potential shortcut artefacts (different columns) without (top three rows) and with post-processing (bottom row). Spoofed data without post-processing are from A17 (TTS), A21 (TTS), and A25 (VC). The bottom row shows the distribution from all the evaluation data after post-processing.

### v4

@@ Abstract
ASVspoof 5 is the fifth edition in a series of challenges which promote the study of speech spoofing and deepfake attacks as well as the design of detection solutions. We introduce the ASVspoof 5 database which is generated in a crowdsourced fashion from data collected in diverse acoustic conditions (cf. studio-quality data for earlier ASVspoof databases) and from $\sim$2,000 speakers (cf. $\sim$100 earlier). The database contains attacks generated with 32 different algorithms, also crowdsourced, and optimised to varying degrees using new surrogate detection models. Among them are attacks generated with a mix of legacy and contemporary text-to-speech synthesis and voice conversion models, in addition to adversarial attacks which are incorporated for the first time. ASVspoof 5 protocols comprise seven speaker-disjoint partitions. They include two distinct partitions for the training of different sets of attack models, two more for the development and evaluation of surrogate detection models, and then three additional partitions which comprise the ASVspoof 5 training, development and evaluation sets. An auxiliary set of data collected from an additional 30k speakers can also be used to train speaker encoders for the implementation of attack algorithms. Also described herein is an experimental validation of the new ASVspoof 5 database using a set of automatic speaker verification and spoof/deepfake baseline detectors. With the exception of protocols and tools for the generation of spoofed/deepfake speech, the resources described in this paper, already used by participants of the ASVspoof 5 challenge in 2024, are now all freely available to the community.
@@ 1 Introduction
The potential threats and risks related to the misuse of deepfakes—synthetic or manipulated media generated with the aid of deep learning—are well acknowledged both by various professional communities (e.g. security, forensics, biometrics) and government bodies, as well as by the general public. Having emerged as a novel type of information security threat which covers multiple domains—news, social media, and communication to name a few—it is more timely than ever to develop proactive defences against deepfakes. The focus of this article is specifically speech deepfakes. Speech, as the primary means of human communication, carries not only the linguistic message but also paralinguistic (beyond language) information. This includes voice timbre in addition to other personally identifiable attributes, representations of which can be estimated from just a few seconds of speech and then used to infer the speaker identity using automatic speaker verification (ASV) technology. The voice timbre can then be transplanted into synthetic or manipulated speech to impersonate a specific, target individual.
@@ 2.2 ASVspoof 5 database partitions
Figure 2: An illustration of the MLS English source database partitioning scheme and its use in generation of the ASVspoof 5 database, including training, development and evaluation sets illustrated in grey boxes. Numbers in the second row indicate the number of speakers in each partition. Those in rectangular blocks denote the number of utterances. The green box denotes the 8 TTS/VC systems used for generation of the ASVspoof 5 training set. The blue box denotes 8 TTS/VC systems used for generation of the development set whereas the red box denotes the 16 TTS/VC systems and adversarial attacks used for generation of the evaluation set. Solid black and coloured arrows denote use of data for the training of TTS/VC attacks or surrogate models. Dashed arrows denote the adaptation towards target speaker and attack generation using input text or source speech utterance.
@@ 3.1 Attack types
TTS-based attacks are performed using a synthetic utterance which contains the voice of a target speaker learned from one or more adaptation utterances and the linguistic content of a provided text input. As illustrated in Figure 4, TTS systems follow one of two general approaches, namely zero-shot voice cloning or few-shot speaker adaptive TTS.121212 Some legacy approaches such as unit-selection TTS are neither zero-shot nor few-shot in strict terms. To help clarify the concept, we consider the operation of the unit-selection TTS, which is used in the ASVspoof 5 database, as a special form of few-shot speaker-adaptive TTS. These systems synthesize speech by selecting and concatenating small speech segments called units. The acoustic decoders (i.e., the selection algorithms) extract units from the adaptation utterance(s). The resulting sequence of units is then simply concatenated by the ‘vocoder’. Readers are encouraged to refer to TTS literature for an in-depth explanation of legacy approaches [37]. A text encoder is used to transform the input text into features which (explicitly or implicitly) encode word pronunciation information and other suprasegmental attributes (e.g., pitch and rhythm). The acoustic decoder then produces acoustic features (e.g., mel-spectrograms), which are further transformed by the vocoder into a speech waveform. [^13] The modules are not necessarily implemented in the form of DNNs. They can instead be a set of linguistic rules or decision trees. So-called end-to-end TTS systems in the form of a single DNN can be divided conceptually into distinct blocks, each of which corresponds to one of the modules illustrated in Figure 4. The two approaches illustrated in Figure 4 differ in how the target speaker voice is learned. In the case of zero-shot voice cloning (top-middle), a target speaker embedding is extracted from the adaptation utterances by a speaker encoder. The acoustic decoder, and optionally also the text encoder and vocoder, are then conditioned upon the target speaker embedding. Since they use target speaker embeddings, zero-shot voice cloning TTS systems are pre-trained using data collected from a large number of speakers, but are not fine-tuned at the level of the target speaker. This sets zero-shot voice cloning systems apart from few-shot speaker adaptive TTS systems (bottom-middle) which are fine-tuned to the voice of the target speaker using the set of adaptation utterances and their transcriptions.
@@
VC-based attacks are implemented using an input utterance containing the voice of a non-target speaker in place of input text, but again using one or more adaptation utterances which contain the voice of a target speaker. The voice in the input utterance is converted to that of the target speaker, whereas the linguistic content of the former is preserved. Approaches to VC can be categorised similarly to the TTS systems illustrated in Figure 4, though with input utterances in place of input text and with a waveform encoder in place of the text encoder. VC acoustic decoders typically use similar statistical models or deep neural networks (DNNs) as TTS, but with necessary modifications to accommodate input features of different dimension.
@@
A28 Zero-shot TTS TF encoder Normalizing flow H/ASP HiFi-GAN YourTTS, pre-trained
A29 Zero-shot TTS GPT-2 - H/ASP HiFi-GAN XTTS [55], pre-trained
@@ 3.2 Training set
A01: a neural TTS system based on Glow-TTS [43]. At the core is an acoustic decoder which generates mel-spectrograms via a normalising-flow-based deep generative model named Glow [56]. The acoustic decoder is conditioned on latent linguistic features which are extracted from the input text by a text encoder using the Transformer encoder architecture [57]. The vocoder is a generative adversarial network (GAN)-based DNN named HiFi-GAN [58]. The speaker encoder is an ECAPA-TDNN [59]. Glow-TTS is open-sourced and is widely used within the TTS community.
@@
A04: a neural TTS system based on Grad-TTS [44], featuring a diffusion-based acoustic decoder. The acoustic decoder generates a mel-spectrogram via an iterative reverse diffusion process [62]. From latent linguistic features generated by the text encoder, the acoustic decoder with a U-Net architecture [63] is used to predict a ‘residual’ (or, more specifically, the gradient field which maximises the probability of data to be generated). The residual is then summed with the latent features, the output of which is again fed to the acoustic decoder to produce another residual. The process is repeated ten times, and the output of the last iteration is used as a generated mel-spectrogram. The text encoder is a Transformer encoder with text and speaker embedding inputs. The vocoder is a HiFi-GAN. The implementation is based on the code released by the original paper, but speaker embeddings are extracted using an ECAPA-TDNN speaker encoder.
@@
A08: a variational inference with adversarial learning for end-to-end text-to-Speech (VITS) [46] system which uses a Transformer-based text encoder and a HiFi-GAN vocoder. The acoustic decoder is based on normalising flow, similar to A01-A03. However, the text encoder, the acoustic decoder, and the vocoder are jointly optimised by maximizing a variational evidence lower bound [46]. Speaker embeddings are x-vectors [64]. VITS produces especially high quality synthetic speech and is readily accessible with the open-sourced toolkit ESPNet [65].
@@ 3.3 Development set
A11: a zero-shot TTS system based on Tacotron 2 [48]. The speaker encoder is a three-layer recurrent neural network (RNN) with long-short-term-memory (LSTM) units [71], trained for a speaker diarization task using generalised end-to-end (GE2E) loss [72]. The text encoder comprises convolution and recurrent layers. The acoustic decoder uses a location-sensitive attention block [73] and an autoregressive (AR) DNN [48] to generate a mel-spectrogram which is then converted into a speech waveform using the diffusion-based WaveGrad [74] vocoder. Tacotron 2 has been reported to produce speech of near-to-natural quality for single-speaker, neutral style TTS [48].
@@
A13: a zero-shot VC (ZSVC) system based on the StarGAN-ZSVC model [49]. The waveform encoder extracts a mel-spectrogram from the input utterance. The StarGAN [75] acoustic decoder then converts the mel-spectrogram so that it encodes the same linguistic content but the voice of the target speaker. The speaker embedding is extracted using a stack of recurrent layers using gated recurrent units (GRUs) [76]. The vocoder is WaveGlow [77], a normalising-flow-based waveform generation model. StarGANs are known to perform well in voice conversion tasks  [78] and the StarGAN-ZSVC model is available as open-source [49].
A14: a zero-shot neural TTS system built upon YourTTS [50], which is itself based on VITS [46]. A14 is a variant of A08. Advances include the higher number of Transformer blocks used in the text encoder and use of a RawNet3 speaker encoder [79].
@@ 3.4 Evaluation set
A24: a zero-shot VC system. An RNN-CNN waveform encoder is used to extract a phonetic posteriorgram [97, 83], a latent representation of mostly speaker-independent phonetic information. The acoustic decoder, also a CNN, combines the posteriorgram with ECAPA-TDNN speaker embeddings to predict target-speaker-dependent latent features. A HiFi-GAN vocoder is then used to produce speech in the voice of the target speaker. All components are trained in an end-to-end manner. A24 is similar to A16 in that input speech is disentangled into speaker-dependent and speaker-independent features. A key difference to A16 is use of GAN-based, end-to-end training.
@@
A29: another pre-trained, zero-shot TTS system, named XTTS [55]. Designed for multi-lingual TTS and trained using a mix of four datasets containing speech in 16 languages, it uses the GPT-2 text encoder architecture [101], a HiFi-GAN vocoder, and an H/ASP speaker encoder (same as for A28). The XTTS system is also publicly available in the Coqui toolkit.
@@ 4.1 Encoding and compression
Condition C00 serves as a reference for which there is no encoding or compression. For conditions C00-C07, all data is sampled at a rate of 16 kHz. In contrast, conditions C08-C11 correspond to a narrow band setting under which all data is down-sampled to 8 kHz. Conditions C01-C11 all involve the application of lossy encoding/compression algorithms at the bitrates indicated in Table 3. C04 and C07 use the recent, deep-learning-based Encodec [102] scheme. For C07, mp3 encoding is applied prior to Encodec to simulate a channel with double compression. Finally, C11 is an experimental setting in which utterances are linearly convolved with the pre-computed, non-linear responses of a set of eight end-to-end calling pipelines, namely calls made from a device to a call center platform over a public switched telephone network (PSTN). Responses were estimated using the synchronised swept sine approach described in [103] using one of six different calling devices and one of four different audio injection methods. Specific configurations for C11 are detailed in A.
@@ 4.2 Shortcut artefacts
We created a post-processing pipeline to detect and reduce the impact of potential shortcut artefacts. It was applied to the ASVspoof 5 evaluation and development sets, but not to the training set. This choice encourages database users to explore techniques which are able to extract the most semantically relevant cues while also avoiding the pitfalls of shortcut learning. Our analysis revealed five potential shortcut artefacts: the peak waveform amplitude; the durations of the leading and trailing non-speech segments [13, 14]; the duration and energy of the whole utterance. The non-speech and speech segments of the utterance in the ASVspoof 5 evaluation set are annotated using the Whisper system [105]. Figure 5 shows the distribution of each shortcut artefact for bona fide utterances (black profiles) and spoofed/deepfake utterances (grey profiles) before post-processing. Distributions are shown in the first three rows for utterances generated using A17 (TTS), A21 (TTS), and A25 (VC) attacks. They show that, for A17 and A25 attacks, the peak amplitude is rescaled to 1.0, a distinct difference to the distribution for bona fide utterances. Whereas there is little difference between the distributions of leading non-speech segment durations for bona fide and spoofed/deepfake utterances for A25, they tend to be of shorter duration for A21. For A17, the duration of leading non-speech segments is uniformly distributed, even if the range is similar to that of bona fide utterances. Even so, the range in the duration of trailing non-speech segments for A17 differs to that for bona fide utterances. The duration of trailing non-speech segments for A21 is generally again shorter than those for bona fide utterances. The distribution of total duration for A25 overlaps with that of bona fide data. This is because A25, a VC system, does not alter the total duration of the input utterance. In the case of TTS-based attacks A17 and A21, the total duration is controlled by the acoustic model, hence the difference between distributions for bona fide and spoofed/deepfake utterances in these cases. There are also distinct differences in distributions of average energy for all three attacks.
@@
Figure 5: Distributions of potential shortcut artefacts (different columns) without (top three rows) and with post-processing (bottom row). Spoofed data without post-processing are from A17 (TTS), A21 (TTS), and A25 (VC). The bottom row shows the distribution from all the evaluation data after post-processing. The equal error rate (EER) shown at the top of each sub-figure is computed using the shortcut artefacts of the bona fide and spoofed utterances. An EER closer to 50% suggests a greater overlap between the distributions of the two classes.